Fermi National Laboratory


Accelerator Update for October 14, 2005 - October 17, 2005

Friday October 14
The day shift began with Operations monitoring store 4445 and a Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beam, and with NuMI, MiniBooNE, and the Meson MCenter experiments taking beam.

At 9 AM, a Linac RF station (LRF3) tripped off and wouldn't reset. An expert began work on the station.

At 9:36 AM, LRF2 tripped off twice. It reset both times.

Main Injector (MI) experts began a study period at 10:29 AM.

At 10:44 AM, Operations reported that the MI PSLOOP beam permit dropped on a door fault. All the VCBs opened. Operators investigated and believed that a MI-10 door was nudged open. Everything eventually reset.

A NuMI expert reported at 11:14 AM that they were turning off beam for Horn chiller work.

At 12:25 PM, LRF3 tripped off for the first of three times in fifteen minutes. It reset each time.

Pbar experts began a reverse proton study at 1:19 PM.

A LCW expert reported at 2:55 PM that the Meson MS2 LCW system had lost 6-8 inches of water since last checked on Thursday. An operator contacted the Switchyard (SY) Coordinator. Operators turned off the system at 3:16 PM and prepared the area for an access to search for a leak.

The NuMI Horn chiller work ended at 4:06 PM. Operations reestablished beam to the experiment at 4:15 PM.

Operations reported at 4:53 PM that they had found the Meson leak from a power supply (MT5Q2). A power supply expert investigated and reported that the leak was on the reversing switch and ran down through the supply soaking everything. He put a heater on the supply and said he'd make repairs on Monday.

Operations established beam to MCenter at 7:21 PM.

The Pbar studies ended and Pbar resumed stacking at 8:24 PM.

Operations began preparing the Recycler to stash antiprotons at 8:48 PM. The stashing ended and Pbar resumed stacking. MCenter resumed taking beam.

Saturday October 15
The midnight shift began with Operations monitoring store 4445 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations terminated store 4445 at 2 AM and then put the Tevatron (TeV )into shot setup.

A Booster RF station (BRF12) tripped off at 2:08 AM for the first of a bunch of times over the next two hours. It reset each time.

At 2:22 AM, there was a smoke alarm at the TeV sector C1 service building. The C1 LCW system also tripped off. Operators found smoke in the building and water coming from the C1 pump. The Duty Electrician cut power to the pump. LCW experts were called in to start up the C3 pump. Then technicians began replacing the C1 pump.

At 3:26 AM, NuMI tripped off for the first of several times over a couple of hours due to a low-pressure indication. It reset each time.

Operations established a Recycler only store 4446 at 5:55 AM with an initial luminosity of 103.8E30.

At 6:45 AM, Operations reported that they had been having trouble with a Booster correction element power supply (B:CEPSW1). Operations called in a Booster expert. At some point after the repair there was a Recycler stash.

At 8:06 AM, a vacuum pump (IP9144) tripped off in the Pbar transfer line A150. Operators tried to shut valves off, but the vacuum problem cascaded to the TeV F sector and caused the store to abort. It was a clean abort with no quench.

Operations reported at 8:27 AM that the F0 vacuum had recovered.

Operations reported by 9:50 AM that they had reestablished beam to NuMI, Pbar, and MiniBooNE.

TeV experts began a study period at 10:34 AM.

At 10:39 AM, Operations reported that BRF12 had tripped off a number of times in the last twenty minutes. Eventually, operators bypassed the station. At 11:51 AM a RF expert brought the station back on line.

At 1 PM, BRF12 tripped off. It reset okay.

The Recycler began preparing at 1:38 PM to stash.

Operations put the TeV into a dry squeeze at 1:48 PM.

The stash ended. Unfortunately, part of the last stash was lost. Pbar resumed stacking at 2:26 PM.

MCenter and MTest resumed taking beam at 2:45 PM.

Operations put the TeV into shot setup at 2:59 PM.

Operations established Recycler only store 4448 at 5:24 PM with an initial luminosity of 86.49E30. Pbar resumed stacking. MCenter and MTest resumed taking beam.

At 7:25 PM, part of the Recycler's Electron Cooling system tripped off. Per an expert's instruction, an operator reset part of it, but left another part off.

Recycler began preparing at 8:26 PM to stash. The stash ended at 9:10 PM. Pbar, MCenter, and MTest resumed normal operations.

Sunday October 16
The midnight shift began with Operations monitoring store 4448 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, MTest, and MCenter taking beam.

MTest reported at 1:23 AM that they were through with the beam for the night.

The Recycler began preparing at 2:38 AM to stash. Operations reported a problem with injections from Pbar into MI. Operators investigated and called in experts. Experts determined that MI wasn't the problem and the stash was completed. At 4:15 AM, Pbar resumed stacking.

At 4:35 AM, Operations reported that a problem with the I:BEAM readback was causing a QXR problem and had also caused the MI injection problem. Operators were investigating both problems. The QXR problems halted all SY beam

Instrumentation technicians came in at 6:21 AM to repair I:BEAM. Operations reported at 8:07 AM that I:BEAM was operating normally.

The Recycler began preparing to stash around 8:14 AM. The stash ended and Pbar and MCenter resumed normal operations.

At 11:05 PM, BRF16 tripped off. It reset okay.

At 12:29 PM, BRF14 tripped off. It would not reset. Operations contacted an expert who said that the cavity strainer needed cleaning and that would require a Booster tunnel access. BRF14 was bypassed

The Recycler began preparing at 3:06 PM to stash. The stash ended and Pbar resumed stacking at 3:44 PM. MCenter resumed taking beam.

Operations terminated store 4448 at 5:03 PM

Operations put the TeV into shot setup at 5:15 PM.

Operations established Recycler only store 4449 at 7:31 PM with an initial luminosity of 126.16E30. Pbar and MCenter resumed normal operations.

At 8:37 PM, BRF12 tripped off. It reset okay.

At 9:39 PM, Operations reported that the Electron Cooling, though left off, had tripped off.

The Recycler began preparing at 9:45 PM to stash. The stash ended and Pbar and MCenter resumed normal operations at 10:54 PM.

Monday October 17
The midnight shift began with Operations monitoring store 4449 and a small stash, with Pbar stacking, and with NuMI, MiniBooNE and MCenter taking beam.

At 1:38 AM, BRF15 tripped off. It reset okay.

The Recycler began preparing at 3:15 AM to stash. The stash ended and Pbar and MCenter resumed normal operations at 4:21 AM.

At 4 AM, BRF15 tripped off. It reset okay.

At 5:17 AM, the NuMI Horn tripped off. It reset okay.

Operations turned off NuMI beam at 7:41 AM for maintenance on a bad chiller compressor.

The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator is planning an eight-hour access into F sector later this week to find and fix a vacuum problem.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 10/17/2005   email Fermilab